OpenStudy (anonymous):

3x – 2y = -1 and 2x + 3y = 3 Determine whether the following lines are parallel, perpendicular, or neither

OpenStudy (anonymous):

need answer

OpenStudy (callisto):

What is the slope of the line having the equation 3x – 2y = -1?

OpenStudy (anonymous):

3

OpenStudy (callisto):

Nope. Either arrange the equation in the form y=mx+c, where m = slope or use slope = -A/B of the equation Ax+By=C to get the slope.